OpenStudy (j2lie):

what mulitplies to -20 and adds up to -22

OpenStudy (anonymous):

XY = -20 X + Y = -22

OpenStudy (j2lie):

that's not what I meant.

OpenStudy (anonymous):

Then what did you mean?

OpenStudy (j2lie):

the factors of 20 that equals to 22

OpenStudy (agent0smith):

List all the factors of 20 20 1 and 20 2 and 10 4 and 5 so, there's none.

OpenStudy (anonymous):

Oh right.

OpenStudy (j2lie):

I realized that so I did something wrong

OpenStudy (j2lie):

I don't know how to factor 12x2-22x-20

OpenStudy (agent0smith):

I see your mistake... that method only works when the a=1 in the ax^2 Factor out the common factor of 2 from everything first... 12x^2-22x-20 = 2(6x^2 - 11x - 10) now you need to multiply the 6 and -10 to get -60 Now find factors of -60 that add up to -11

OpenStudy (j2lie):

you're so smart.

OpenStudy (agent0smith):

Haha thank you :)

OpenStudy (j2lie):

I forgot about that. The sub teacher is not good at math.

OpenStudy (agent0smith):

hehe yeah it's not often you get a sub who has a degree in math and physics :D

OpenStudy (j2lie):

the sub does it really fast.

OpenStudy (j2lie):

-6 + 1 does add up to -11.

OpenStudy (j2lie):

actually you're suppose to multiply the 6 and the 10.

OpenStudy (agent0smith):

2(6x^2 - 11x - 10) -60 -15 and 4 which add to -11 So break up the -11x into -15x and 4x: 2(6x^2 - 15x + 4x - 10) Now try to factor by grouping

OpenStudy (triciaal):

thanks Agent0smith i was stuck because i did not stop to factor i automatically was working with the numbers that popped out

OpenStudy (agent0smith):

2(6x^2 - 15x + 4x - 10) factor by grouping 2( 3x(2x - 5) + 2(2x - 5) ) then finally... 2(2x - 5)(3x + 2)

OpenStudy (j2lie):

And I told you that.

OpenStudy (agent0smith):

told us what...?

OpenStudy (j2lie):

I was trying to tell her your way she wasn't listening.